# A Characterization for Tightness of the Sparse Moment-SOS Hierarchy

### v1

@@ 3.1. Characterization for the tightness
First, we give a sufficient and necessary condition for the sparse Moment-SOS hierarchy of (3.2)-(3.3) to be tight. The following is the main theorem.
@@ 3.2. Detecting tightness and extracting minimizers
Proof.
Since (3.8) holds, we have
$\rho\coloneqq\min_{1\leq i\leq m}\rho_{i}.$
For the case $10<\rho<$ 1, define the new tms
$\mathscr{V}_{h_{i}}^{spa,2t}[y^{*}]$ $=$ $\rho\mathscr{V}_{h_{i}}^{spa,2t}[\hat{y}]+(1-\rho)\mathscr{V}_{h_{i}}^{spa,2t}[\tilde{y}],$
$\mathscr{V}_{h_{i}}^{spa,2t}[\tilde{y}]$ $=$ $\frac{1}{1-\rho}\left(\mathscr{V}_{h_{i}}^{spa,2t}[y^{*}]-\rho\mathscr{V}_{h_{i}}^{spa,2t}[\hat{y}]\right)=0,$
On the other hand, it also holds that
Since $10<\rho<$ 1, we must have
$f(x^{*})=f_{t}^{smo}\leq f_{\min}\leq f(x^{*}).$
This shows that $x^{*}$ is a minimizer of (1.1). ∎
@@
When (3.10) holds, we have the decomposition (3.8) with $r_{i}=\mbox{rank}\,M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]$. We refer to [6, 18] for this fact. Under some genericity assumptions, the dense Moment-SOS hierarchy is tight if and only if the flat truncation holds (see [27, Theorem 2.2] or [30, Section 5.3]).
@@
Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$. Assume $y^{*}$ is a minimizer of (3.3) and (3.10) holds for each $i$. Let $r_{i}$ be as in (3.8). Assume $r_{1}=\cdots=r_{m}=r$. If for all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$ it holds
@@
Proof.
Pick arbitrary $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, then (3.12) holds, by the given assumption. The flat truncation (3.10) implies the decomposition (3.8). Note that
Up to permutation, we have
Since $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$, the above implies that $\lambda_{i,l}=\lambda_{j,l}$ for all $i,j,l$. Moreover, all $u^{(i,l)},u^{(j,l)}$ are projections of a common vector in $\mathbb{R}^{\Delta_{i}\cup\Delta_{j}}$. Therefore, there are points $u^{(1)},\ldots,u^{(r)}\in\mathbb{R}^{n}$ such that (3.13) holds. Since $(u^{(j)})_{\Delta_{i}}\in K_{\Delta_{i}}$ for all $i$, we have $u^{(j)}\in K$ for all $j$. So every $[u^{(j)}]_{spa,2t}$ is feasible for (3.3) with the relaxation order equal to $t$, and every $[u^{(j)}]_{spa,2k}$ is feasible for (3.3) with the relaxation order equal to $k$. Therefore,
for all $j=1,\ldots,r$. Moreover, the decomposition (3.13) implies
Since each $\lambda_{i}>0$, the above implies that for all $j=1,\ldots,r$,
We remark that if the set $\mathcal{D}=\{\Delta_{1},\ldots,\Delta_{m}\}$ is not a connected cover of $[n]$, then $\mathcal{D}$ can be expressed as a union of disjoint subsets, say, $\mathcal{D}=\mathcal{D}_{1}\cup\cdots\cup\mathcal{D}_{\ell},$ such that each $\mathcal{D}_{i}$ is a connected cover for a subset of $[n]$. Then, the sparse optimization problem (1.1) can be split into a union of smaller sized problems that do not have mutually joint variables.
@@ 4. Some sufficient conditions for tightness
We remark that (4.1) does not imply $f_{i}+p_{i}\in\mbox{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$. This section explores various conditions for Assumption 4.1 to hold and for the sparse Moment-SOS hierarchy of (3.2)-(3.3) to be tight.
@@ 4.1. The convex case
Assume each $h_{i}$ is linear (or it does not appear) and $u$ is a minimizer of (1.1). Under certain constraint qualification like Slater’s condition (i.e., there is a feasible point $x$ for (1.1) such that all $g_{i}(x)>0$), there exist Lagrange multipliers $\lambda_{i,j}$ and $\nu_{i,j}$ such that ($\perp$ means that the product is 00)
@@
This is called the first order optimality condition (FOOC).
@@
$\nabla_{x_{\Delta_{i}}}^{2}\bigg{[}f_{i}(u_{\Delta_{i}})-\sum_{j=1}^{s_{i}}\lambda_{i,j}g_{i,j}(u_{\Delta_{i}})\bigg{]}$
@@
(4.4) $\left\{\begin{array}[]{cl}\displaystyle\min_{x_{\Delta_{i}}}&f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})\\ \mbox{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0.\end{array}\right.$
Since $p_{i}(x_{\Delta_{i}})$ is linear, $f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})$ is also convex in $x_{\Delta_{i}}$, hence $u_{\Delta_{i}}$ is a minimizer of (4.4). Its minimum value is 00, so $f_{i}+p_{i}\geq 0$ on $K_{\Delta_{i}}$. Therefore, Assumption 4.1 holds. The tightness of the hierarchy of (3.2)-(3.3) follows from Theorem 3.1 and [4, Corollary 3.3] (or [30, Theorem 7.2.5]). ∎
Theorems 3.3 and 3.4 can be applied to get minimizers of (1.1). Note that a convex optimization problem typically has a unique optimizer, or otherwise it has infinitely many optimizers. For this reason, when the conditions (3.8)-(3.9) and/or (3.12)-(3.13) hold, we typically have $r_{i}=1$ for all $i$. For such cases, if $y^{*}$ is a minimizer of (3.3), the point
@@
Next, we consider the special case that the defining polynomials of (1.1) are SOS-convex. Recall that a polynomial $\phi(x)$ is SOS-convex if its Hessian $\nabla^{2}\phi(x)=H(x)^{T}H(x)$ for some matrix polynomial $H(x)$. We refer to [30, Chap. 7] for SOS-convex polynomials. Under the SOS-convexity assumption, the sparse Moment-SOS relaxations (3.2)-(3.3) are tight for all $k\geq k_{0}$ and the condition (3.5) holds.
@@
Suppose $y$ is feasible for the relaxation (3.3) and let $v=(y_{e_{1}},\ldots,y_{e_{n}})$. Since they are SOS-convex, by Jensen’s inequality (see [16] or [30, Chap. 7]), we have
@@ 4.2. The case of sufficient optimality conditions
Under the archimedianness, the dense Moment-SOS hierarchy in [14] is tight when linear independence constraint qualification (LICQC), strict complementarity condition (SCC) and second order sufficient condition (SOSC) hold at every minimizer, as shown in [29]. We refer to [30, Chapter 5.1] for these conditions. For sparse Moment-SOS relaxations, we have a similar conclusion under Assumption 4.1.
@@
(4.6) $\left\{\begin{array}[]{rl}\min&f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})\\ \mbox{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0.\end{array}\right.$
@@
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.4). Assume (4.8) holds for all $i$ and let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
@@
Proof.
(i) Since each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$ is archimedean, the feasible sets $K_{\Delta_{i}}$ and $K$ are all compact. So, (1.1) achieves its minimum value, say, at a minimizer $u\in K$. Assumption 4.1 implies
$(f_{1}+p_{1})+\cdots+(f_{m}+p_{m})=f-f_{\min}.$
Since $f(u)=f_{\min}$ and each $f_{i}+p_{i}\geq 0$ on $K_{\Delta_{i}}$, the minimum value of (4.6) is 00. By [29, Theorem 1.1], if the LICQC, SCC, and SOSC hold at every minimizer of (4.6), then there is a relaxation order $N_{i}$ such that for all $\epsilon>0$,
Let $N\coloneqq\max\{N_{1},\ldots,N_{m}\}$. The above then implies
By Assumption 4.1, we have $p_{1}+\dots+p_{m}+f_{\min}=0$, so for all $k\geq N$
$\in$ $\operatorname{Ideal}[h]_{spa,k}+\operatorname{QM}[g]_{spa,k}.$
This follows from [23, Theorem 9.5.3] and [29, Theorem 3.1]. So, we get
By Assumption 4.1, $p_{1}+\cdots+p_{m}+f_{\min}=0$, so
(4.13) $\left\{\begin{array}[]{cll}\max&\gamma\\ \mbox{s.t.}&f_{i}+p_{i}-\gamma\in\operatorname{Ideal}_{\Delta_{i}}[h_{i}]_{2k}+\operatorname{QM}_{\Delta_{i}}[g_{i}]_{2k}.\end{array}\right.$
Its dual optimization is the moment relaxation
(4.14) $\left\{\begin{array}[]{cll}\min&\langle f_{i}+p_{i},y_{\Delta_{i}}\rangle\\ \mbox{s.t.}&\mathscr{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]=0,L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}]\succeq 0,\\ &(y_{\Delta_{i}})_{0}=1,\ M_{\Delta_{i}}^{(k)}[y_{\Delta_{i}}]\succeq 0,\\ &y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}^{\Delta_{i}}_{2k}}.\end{array}\right.$

### v2

@@ 3.1. Characterization for the tightness
First, we give a sufficient and necessary condition for the sparse SOS hierarchy of (3.2) to be tight. We remark that the sparse Moment-SOS hierarchy of (3.2)-(3.3) is tight if and only if the sparse SOS hierarchy of (3.2) is tight. The following is the main theorem.
@@ 3.2. Detecting tightness and extracting minimizers
The proof of Theorem 3.3 is given in Subsection 7.1. For the dense Moment-SOS hierarchy (i.e., $m=1$), the tightness is guaranteed if there exists a degree $t\in[k_{0},k]$ such that (3.8) holds. This is because for the dense case, we have (let $r\coloneqq r_{i}$, $u^{(j)}\coloneqq u^{(1,j)}$ and $\lambda^{(j)}\coloneqq\lambda^{(1,j)}$ for notational convenience)
$f(u^{(j)})\geq f_{\min}\geq\left<f,y^{*}|_{2k}\right>=\left<f,y^{*}|_{2t}\right>=\lambda_{1}f(u^{(1)})+\cdots+\lambda_{r}f(u^{(r)}),$
which implies
$f(u^{(1)})=\cdots=f(u^{(r)})=\left<f,y^{*}|_{2k}\right>=f_{\min}.$
However, for the sparse Moment-SOS hierarchy of (3.2)-(3.3), even if (3.8) holds for all $i$, we may not have $f^{smo}_{k}=f_{\min}$ and there may not exist a point $x^{*}=(x_{1}^{*},\,\cdots,\,x_{n}^{*})$ such that every $x_{i}^{*}\in\mathcal{X}_{\Delta_{i}}$. This is demonstrated as follows.
Example 3.4.
Consider the following sparse polynomial optimization problem:
(3.10) $\left\{\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{3}}&f(x)=x_{1}+x_{2}+x_{3}\\ \mathrm{s.t.}&(x_{1}-1)(x_{1}-2)(x_{1}+x_{2}-6)=0,\ x_{1}-x_{2}=0,\\ &(x_{2}-1)(x_{2}-2)(x_{2}+x_{3}-6)=0,\ (x_{2}+x_{3}-3)(x_{2}-3)=0,\\ &(x_{3}-1)(x_{3}-2)(x_{1}+x_{3}-6)=0,\ x_{1}-x_{3}=0.\\ \end{array}\right.$
In the above, $f_{1}=x_{1},f_{2}=x_{2},f_{3}=x_{3},$ $\Delta_{1}=\{1,2\}$, $\Delta_{2}=\{2,3\}$, $\Delta_{3}=\{1,3\}$, and
$h_{1}=\big{(}(x_{1}-1)(x_{1}-2)(x_{1}+x_{2}-6),\,x_{1}-x_{2}\big{)},$
$h_{2}=\big{(}(x_{2}-1)(x_{2}-2)(x_{2}+x_{3}-6),\,(x_{2}+x_{3}-3)(x_{2}-3)\big{)},$
$h_{3}=\big{(}(x_{3}-1)(x_{3}-2)(x_{1}+x_{3}-6),\,x_{1}-x_{3}\big{)}.$
Clearly, we can check that $K=\{(3,3,3)\}$ and
(3.11) $K_{\Delta_{1}}=\Big{\{}\begin{bmatrix}1\\ 1\end{bmatrix},\begin{bmatrix}2\\ 2\end{bmatrix},\begin{bmatrix}3\\ 3\end{bmatrix}\Big{\}},$ $K_{\Delta_{2}}=\Big{\{}\begin{bmatrix}1\\ 2\end{bmatrix},\begin{bmatrix}2\\ 1\end{bmatrix},\begin{bmatrix}3\\ 3\end{bmatrix}\Big{\}},$
$K_{\Delta_{3}}=\Big{\{}\begin{bmatrix}1\\ 1\end{bmatrix},\begin{bmatrix}2\\ 2\end{bmatrix},\begin{bmatrix}3\\ 3\end{bmatrix}\Big{\}}.$
For each $k\geq 2$, let $y^{*}\in\mathbb{R}^{\mathbb{U}_{k}}$ be such that
(3.12) $y^{*}_{\Delta_{1}}=\frac{1}{2}\Big{(}\begin{bmatrix}1\\ 1\end{bmatrix}_{2k}+\begin{bmatrix}2\\ 2\end{bmatrix}_{2k}\Big{)},$ $y^{*}_{\Delta_{2}}=\frac{1}{2}\Big{(}\begin{bmatrix}1\\ 2\end{bmatrix}_{2k}+\begin{bmatrix}2\\ 1\end{bmatrix}_{2k}\Big{)},$
$y^{*}_{\Delta_{3}}=\frac{1}{2}\Big{(}\begin{bmatrix}1\\ 1\end{bmatrix}_{2k}+\begin{bmatrix}2\\ 2\end{bmatrix}_{2k}\Big{)}.$
The above $y^{*}$ is well-defined, i.e., $(y_{\Delta_{i}}^{*})_{\alpha}=(y_{\Delta_{j}}^{*})_{\alpha}$ for all $\alpha\in\mathbb{N}_{2k}^{\Delta_{i}}\cap\mathbb{N}_{2k}^{\Delta_{j}}$, and it is feasible for the sparse moment relaxtion (3.3). So, $f^{smo}_{k}\leq\langle f,y^{*}\rangle=\frac{9}{2}$. Since
$x_{1}+x_{2}+x_{3}-\frac{9}{2}$ $=$ $(x_{2}-2x_{1}+2)h_{1,2}-h_{2,2}+(2x_{3}-x_{1}-1)h_{3,2}$
$+\frac{3}{2}(x_{1}-x_{2})^{2}+\frac{1}{2}(x_{2}+x_{3}-3)^{2}+\frac{3}{2}(x_{1}-x_{3})^{2}$
$\in$ $\operatorname{Ideal}[h]_{spa,2}+\Sigma[x_{\Delta_{1}}]_{2}+\Sigma[x_{\Delta_{2}}]_{2}+\Sigma[x_{\Delta_{3}}]_{2},$
we have $f^{smo}_{k}\geq f^{spa}_{k}\geq\frac{9}{2}$ for all $k\geq 2$. Hence, $f^{smo}_{k}=f^{spa}_{k}=\frac{9}{2}$ and $y^{*}$ is a minimizer of (3.3). However, $f^{smo}_{k}<f_{\min}=9$, since the only feasible point of (3.10) is $(3,3,3)$. This shows that condition (3.8) is not sufficient for the sparse Moment-SOS hierachy to be tight.
@@
When (3.13) holds, we have the decomposition (3.8) with $r_{i}=\mbox{rank}\,M_{\Delta_{i}}^{(t)}[y^{*}_{\Delta_{i}}]$. We refer to [7, 20] for this fact. Under some genericity assumptions, the dense Moment-SOS hierarchy is tight if and only if the flat truncation holds (see [29, Theorem 2.2] or [32, Section 5.3]). For the sparse Moment-SOS hierarchy, if (3.13) holds with $t=k$ for all $i$ and $\mbox{rank}\,M_{\Delta_{ij}}^{(k)}[y^{*}_{\Delta_{i}}]=1$ for all $i\neq j$ when $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, then $f^{smo}_{k}=f_{\min}$ and one can extract minimizers for (1.1); see [17, Theorem 3.7]. Moreover, by Theorem 3.3, if (3.13) holds for some $t\leq k$ for all $i$ and $f^{smo}_{t}=f^{smo}_{k}$, then each point $x^{*}$ such that every $x_{\Delta_{i}}^{*}\in\mathcal{X}_{\Delta_{i}}$ minimizes (1.1), and tightness of the sparse Moment relaxation (3.3) is certified when such $x^{*}$ exists.
@@
Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ and satisfies the RIP. Assume $y^{*}$ is a minimizer of (3.3) and (3.13) holds for each $i$. Let $r_{i}$ be as in (3.8). Assume $r_{1}=\cdots=r_{m}=r$. If for all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$ it holds
@@
The proof of Theorem 3.5 is given in Subsection 7.2. For the dense case (i.e., $m=1$), if the flat truncation (3.13) holds, then there exists a unique decomposition (up to permutations) given as in (3.16), and the tightness of the dense moment relaxation can be certified. However, for the sparse case, the condition (3.13) may not guarantee existence of (3.16) or tightness of the sparse moment relaxation (3.3). To see this, consider the problem (3.10) in Example 3.4 and the minimizer $y^{*}\in\mathbb{R}^{\mathbb{U}_{2}}$ as in (3.12). Both conditions (3.13) and (3.15) hold for all $i=1,2,3$ with $t=2$ and $r=2$. But the decomposition (3.16) does not hold and $f^{smo}_{k}<f_{\min}$ for all $k$. The RIP fails for Example 3.4. When RIP holds, conditions (3.13) and (3.15) imply the decomposition (3.16) and tightness of the sparse moment relaxation (3.3).
If the set $\mathcal{D}=\{\Delta_{1},\ldots,\Delta_{m}\}$ is not a connected cover of $[n]$, then $\mathcal{D}$ can be expressed as a union of disjoint subsets, say, $\mathcal{D}=\mathcal{D}_{1}\cup\cdots\cup\mathcal{D}_{\ell},$ such that each $\mathcal{D}_{i}$ is a connected cover for a subset of $[n]$. Then, the sparse optimization problem (1.1) can be split into a union of smaller sized problems that do not have mutually joint variables.
@@ 4. Some sufficient conditions for tightness
For the dense case (i.e., $m=1$), Assumption 4.1 holds automatically for $p_{1}=-f_{\min}$ if $-\infty<f_{\min}<+\infty$ (note that $f_{\min}=+\infty$ if $K=\emptyset$). However, it may not hold for the sparse case. We refer to Example 6.7 for such an exposition. Besides that, we remark that (4.1) does not imply $f_{i}+p_{i}\in\mbox{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$. This section explores various conditions for Assumption 4.1 to hold and for the sparse Moment-SOS hierarchy of (3.2)-(3.3) to be tight.
@@ 4.1. The convex case
Assume each $h_{i}$ is linear (or it does not appear), all $f_{i}$ and $-g_{i,j}$ are convex, and $u$ is a minimizer of (1.1). Under certain constraint qualification like Slater’s condition (i.e., there is a feasible point $x$ for (1.1) such that all $g_{i}(x)>0$), there exist Lagrange multipliers $\lambda_{i,j}$ and $\nu_{i,j}$ such that ($\perp$ means that the product is 00)
@@
This is called the first order optimality condition (FOOC). By Theorem 3.1, if Assumption 4.1 is satisfied with polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that every $f_{i}+p_{i}$ is convex, then the sparse Moment-SOS hierarchy is tight under certain conditions; see [32, Chapter 7.2]. Interestingly, when (1.1) has a minimizer satisfying the FOOC (4.3), such polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ always exist. This is shown as follows.
@@
$\nabla_{\Delta_{i}}^{2}\bigg{[}f_{i}(u_{\Delta_{i}})-\sum_{j=1}^{s_{i}}\lambda_{i,j}g_{i,j}(u_{\Delta_{i}})\bigg{]}$
@@
(4.4) $\left\{\begin{array}[]{cl}\displaystyle\min_{x_{\Delta_{i}}}&f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})\\ \mathrm{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0.\end{array}\right.$
Since $p_{i}(x_{\Delta_{i}})$ is linear, $f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})$ is also convex in $x_{\Delta_{i}}$, hence $u_{\Delta_{i}}$ is a minimizer of (4.4). Its minimum value is 00, so $f_{i}+p_{i}\geq 0$ on $K_{\Delta_{i}}$. Therefore, Assumption 4.1 holds. The tightness of the hierarchy of (3.2)-(3.3) follows from Theorem 3.1 and [5, Corollary 3.3] (or [32, Theorem 7.2.5]). ∎
Theorems 3.3 and 3.5 can be applied to get minimizers of (1.1). Note that a convex optimization problem typically has a unique optimizer, or otherwise it has infinitely many optimizers. For this reason, when the conditions (3.8)-(3.9) and/or (3.15)-(3.16) hold, we typically have $r_{i}=1$ for all $i$. For such cases, if $y^{*}$ is a minimizer of (3.3), the point
@@
Next, we consider the special case that the defining polynomials of (1.1) are SOS-convex. Recall that a polynomial $\phi(x)$ is SOS-convex if its Hessian $\nabla^{2}\phi(x)=H(x)^{T}H(x)$ for some matrix polynomial $H(x)$. We refer to [32, Chap. 7] for SOS-convex polynomials. For the dense case (i.e., $m=1$), when $f$ and all $-g_{i,j}$ are SOS-convex and $h$ is linear (or it does not appear), the dense Moment-SOS relaxations are tight for all $k\geq k_{0}$; see [18] or [32, Section 7.2]. In the following, we show that the sparse Moment-SOS relaxations (3.2)-(3.3) are also tight for all $k\geq k_{0}$ and the condition (3.5) holds, under the SOS-convexity assumption.
@@
Suppose $y$ is feasible for the relaxation (3.3) and let $v=(y_{e_{1}},\ldots,y_{e_{n}})$. Since they are SOS-convex, by Jensen’s inequality (see [18] or [32, Chap. 7]), we have
@@ 4.2. The case of sufficient optimality conditions
Under the archimedeanness, the dense Moment-SOS hierarchy in [16] is tight when linear independence constraint qualification condition (LICQC), strict complementarity condition (SCC) and second order sufficient condition (SOSC) hold at every minimizer, as shown in [31]. We refer to [32, Chapter 5.1] for these conditions. For sparse Moment-SOS relaxations, we have a similar conclusion under Assumption 4.1.
@@
(4.6) $\left\{\begin{array}[]{rl}\min&f_{i}(x_{\Delta_{i}})+p_{i}(x_{\Delta_{i}})\\ \mathrm{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0.\end{array}\right.$
@@
    Suppose $\{\Delta_{1},\ldots,\Delta_{m}\}$ is a connected cover of $[n]$ (see the definition before Theorem 3.5) and satisfies the RIP. Assume (4.8) holds for all $i$ and let $\mathcal{X}_{\Delta_{i}}$ be the set as in (3.9). For all $i\neq j$ with $\Delta_{i}\cap\Delta_{j}\neq\emptyset$, suppose the projection
@@
The proof of Theorem 4.4 is given in Subsection 7.3. For the dense case (i.e., $m=1$), the flat truncation (4.8) is sufficient for the moment relaxation to be tight. However, for the sparse case, the condition (4.8) alone cannot guarantee tightness, because it may not imply (4.10). Example 3.4 is an exposition for this.